# ENFORCE: Exact Nonlinear Constrained Learning with Adaptive-depth Neural Projection

### v2

@@ 5 Experiments and discussion
Effects of constrained learning
(a) $y_{1}$
(b) $y_{2}$
(c) Constraint Satisfaction
(d) Data Scarcity
Figure 3: Prediction comparison between ENFORCE ($\lambda=0.5$, $\epsilon_{T}=10^{-4}$, $\epsilon_{I}=10^{-6}$) and a multilayer perceptron (MLP). ENFORCE enhances the overall accuracy and guarantees satisfaction for highly nonlinear constraints. ENFORCE consistently performs better than a standard MLP even when trained on uniformly sampled fractions of the training dataset.
Training dynamics
To understand the training dynamics of ENFORCE, we analyze the loss curves shown in Fig. 4(a), where the training data loss of ENFORCE is compared to the MLP. Being interested in the effect of hard-constrained learning and to ease the visualization, we do not report here the loss curve of the soft-constrained NN. In this case study, AdaNP contributes to the learning process from the very early iterations (Fig. 4(b), orange line), suggesting that the projection operations positively guide the optimization process. The combination of feasible predictions and minimization of projection displacement drives the learning process toward more optimal outcomes.
(a) Loss NN vs. ENFORCE
(b) ENFORCE Losses and AdaNP
Constrained learning hyperparameters (training)
We systematically analyze the influence of hyperparameters, such as the weighting factor $\lambda$ and the tolerance $\epsilon_{T}$, on the constrained learning process. Fig. 5 shows the influence of the hyperparameters on the accuracy of trained ENFORCE models evaluated on the test set.
The regression accuracy is evidently affected by the choice of the loss weighting factor $\lambda$ (Fig. 5). Remarkably, unlike the challenging task of tuning weighting factors in soft-constrained methods [18], the constrained learning approach proposed here positively impacts accuracy regardless of the specific weighting factor chosen (as shown in Fig. 5, the accuracy of ENFORCE is consistently greater than that of a standard MLP). However, an inappropriate choice of this parameter can result in suboptimal outcomes (e.g., when $\lambda=2$ in Fig. 5). Therefore, careful tuning of this hyperparameter is warranted.
Constrained learning hyperparameters (inference)
During inference, ENFORCE dynamically adapts the depth of AdaNP to ensure an average tolerance below $\epsilon_{I}=10^{-6}$ in this case study. The required depth, however, also depends on the training parameters. Fig. 6(b) illustrates the number of NP layers needed to satisfy the constraint under varying $\lambda$ and $\epsilon_{T}$. The weighting factor is shown to reduce the required number of NP layers by half, with no additional cost during training. This phenomenon can be attributed to the fact that, in the absence of a penalty for projection displacement, the neural network is free to learn a function that, although potentially far from the actual one, results in projections that fall within the vicinity of the ground truth. This approach, however, necessitates multiple projections. In contrast, the penalty term drives the model to learn a function that is sufficiently close to the ground truth, thereby reducing the number of neural projections required. Increasing the value of $\epsilon_{T}$ impacts (positively) the depth of AdaNP at inference time when the displacement penalty factor is set to be small during training. This finding further supports the recommendation of employing shallow AdaNP modules during training, by relaxing the value of $\epsilon_{T}$.
We conclude that the weighting factor $\lambda$ plays an important role by balancing the contribution of the projection displacement error. On the other hand, enforcing strict satisfaction during training at arbitrary small tolerance $\epsilon_{T}$ does not necessarily improve the overall outcome.
Figure 5: Influence of constrained learning hyperparameters on the accuracy of ENFORCE on the test set (note that here we plot the inverse of the mean squared error (MSE)). The weighting factor $\lambda$ favors the learning process if appropriately tuned. Conversely, the training tolerance $\epsilon_{T}$ exhibits a small impact on performance, suggesting it can be set based on available resources. Overall, despite the choice of hyperparameters, ENFORCE is more accurate than an MLP with the same complexity, while also satisfying the underlying nonlinear constraint.
Figure 6: Dynamic evolution of AdaNP during training and inference when different training hyperparameters are chosen. At training time, AdaNP is deeper as a smaller constraints tolerance $\epsilon_{T}$ is chosen.
6 Conclusions and outlook
We developed ENFORCE, a method designed to guarantee neural network prediction satisfies any nonlinear constraint $c(x,y)=0$ within a specified tolerance $\epsilon$, without relying on external solvers or incurring prohibitive computational overhead. The effectiveness of the method is demonstrated in an illustrative case study chosen for its interpretability and ease of visualization. We found that (1) ENFORCE ensures the satisfaction of nonlinear constraints to within a tolerance $\epsilon$ for both average and maximum residuals, (2) the overall accuracy improves from traditional unconstrained MLP, achieving near-perfect performance on the test set, and (3) the constrained learning hyperparameters influence the learning process toward optimal performance; however, our method shows superior performance compared the unconstrained counterpart for a wide range of hyperparamer settings. Additionally, we observed promising performance in data-scarcity conditions (Fig. 3(d)).
The method has the potential to address specific challenges or complement existing approaches, including those based on neural networks and other machine learning models for regression tasks (e.g., Gaussian processes and support vector machines). For instance, hard constraints can be combined with soft-constraint techniques, such as PINNs, to reliably solve PDEs [56]. Additionally, the method could enhance learning performance in partially annotated datasets by inferring missing information through available constraints. Finally, an interesting future direction could involve applying the AdaNP module to GenAI models, guiding the generation process toward domain-compliant samples, such as for synthetic data or image/video generation.
Limitations
In this paragraph. we highlight the main limitation of the proposed method. Firstly, the effectiveness of ENFORCE is highly dependent on the regression capabilities of the neural network backbone. When the model lacks sufficient complexity to achieve accurate predictions, ENFORCE provides limited benefit. Furthermore, the method becomes computationally intensive when applied to systems with a large number of constraints (e.g., more than 1000). This is due to the computational cost of each neural projection layer, which scales as $O(N^{3})$ because of the matrix inversion operation, where $N$ is the number of constraints. Further details are provided in Appendix A.2.
Acknowledgements
@@
Appendix A Appendix
A.1 Mathematical derivations
This section provides the key mathematical derivations underlying the proposed method. These derivations are intended to illustrate the theoretical foundations of the approach and to support some of the discussions presented in the main text.
Closed-form neural projection
@@ A.1 Mathematical derivations
Deviation from Newton’s method
To support the discussion raised in Section 4.1, we show how our method deviates from standard Newton’s method for solving nonlinear KKT conditions. Given a NLP:
@@
If we linearized the system according to the Newton’s iteration, we obtain:

### v3

@@ 5.1 Illustrative function fitting
We report the main results in Table 1. Notably, ENFORCE guarantees constraint satisfaction (below the required tolerance) with minimal computational overhead (+6 ms) while also improving predictive accuracy over standard MLPs, addressing the core challenge of balancing expressivity and feasibility in constrained learning. Extensive discussion about performance, constrained learning dynamics, and hyperparameters is reported in Appendix C. Notably, we observe that, after constrained training, ENFORCE outperforms the MLP even before projection (i.e., using just the backbone at inference).
5.2 Constrained optimization problem
A relevant field in neural network research involves learning approximate solutions to large-scale constrained optimization problems. In this context, NNs serve as fast and inexpensive alternatives to traditional solvers, which are often computationally intensive. Following the benchmarking protocol introduced by Donti et al. [7], we first evaluate ENFORCE on learning the solution of nonconvex optimization problems with linear equality constraints. We then extend the evaluation by replacing the linear equality constraints with nonlinear ones. We compare against alternative methods for learning (or solving) constrained optimization problems, such as MLP, soft constrained MLP, the state-of-the-art DC3 [7], and the deterministic nonlinear programming solver IPOPT [31]. For every model, we use an equivalent fully connected ReLU neural network backbone consisting of 2 hidden layers with 200 neurons. Training is performed with a batch size of 200, a learning rate of $10^{-4}$, and until model convergence (3,500 epochs for DC3; 1,000 for ENFORCE and the other baselines). Every run is repeated 5 times. Like Donti et al., we run the test inference on a single batch of 833 samples, and for the inference time of the optimizer, we assume full parallelization on 833 CPUs (cf. [7] for more details).
5.2.1 Nonconvex problem with linear constraints
We consider the same class of nonconvex optimization problems as in Donti et al. [7], with the focus on equality constraints:
$\min_{y\in\mathbb{R}^{N_{O}}}f_{\text{obj}}(y)=\frac{1}{2}y^{T}Qy+p^{T}\sin{y},\quad\text{s.t.}\quad c^{T}y=x,$ (9)
where $Q\in\mathbb{R}^{N_{O}\times N_{O}}\succeq 0$, $p\in\mathbb{R}^{N_{O}}$, and $c\in\mathbb{R}^{N_{C}\times N_{O}}$ are randomly sampled constant parameters, while $x\in\mathbb{R}^{N_{C}}$ (with $N_{C}$ = $N_{I}$) is the variable parameter across problem instances. $Q$ is a diagonal matrix chosen to be positive semi-definite and $x$ is uniformly sampled in the interval $[-5,5]$. We aim to learn the optimal $y$ given an instance of $x$ in an unsupervised fashion. Rather than using a dataset of solved optimization instances, we minimize the objective in the unsupervised task loss $\ell_{T}=f_{\text{obj}}(y)=\frac{1}{2}y^{T}Qy+p^{T}\sin{y}$.
In Table 2, we report the results on the constrained nonconvex task for 200 variables and 150 linear equality constraints. Given the linear nature of the constraints, ENFORCE consistently guarantees the feasibility for all the test samples. ENFORCE learns a solution that is only $0.47\%$ suboptimal relative to the objective value obtained by IPOPT, while obtaining a 25×25$\times$25 × acceleration. ENFORCE shows faster training convergence when compared to DC3 (cf. Table 4, Appendix C) and an optimal objective gain of $40\%$. This improvement can be attributed to the stability of the projection mapping, in contrast to the null-space completion method. As expected, unconstrained and soft-constrained methods do not guarantee feasibility and may yield predicted optima with lower objective values than those computed by constraint-respecting solvers. However, in constrained optimization, such infeasible solutions are inadmissible, regardless of their objective value. In Appendix C (Table 4), we show how the performance of the methods scales with varying numbers of variables and constraints, including training time.
Table 2: Results on a batch of 833 instances of nonconvex optimization problems with linear equality constraints involving 200 variables and 150 equality constraints. ENFORCE guarantees the feasibility of the solutions, is 25×25$\times$25 × faster than IPOPT, although assuming full parallelization, and learns a $40\%$ better optimum than the state-of-the-art DC3 method. Baseline deep learning-based and soft-constrained methods show significant constraint violations and suboptimal predictions.
Method Obj. value Max eq. Mean eq. Inf. time [s]
IPOPT [31] $-10.64\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.379\pm 0.060$
MLP $-52.99\pm 0.01$ $45.38\pm 0.56$ $9.14\pm 0.02$ $0.001\pm 0.001$
Soft ($\lambda_{c}=5$) $1.28\pm 0.32$ $1.45\pm 0.43$ $0.08\pm 0.00$ $0.001\pm 0.000$
Soft ($\lambda_{c}=1$) $-8.18\pm 0.18$ $1.47\pm 0.41$ $0.09\pm 0.00$ $0.001\pm 0.001$
Soft ($\lambda_{c}=0.1$) $-13.55\pm 0.02$ $2.17\pm 0.12$ $0.35\pm 0.00$ $0.001\pm 0.001$
DC3 [7] $-6.27\pm 0.07$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.004\pm 0.000$
ENFORCE $-10.59\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.016\pm 0.002$
5.2.2 Nonconvex problem with nonlinear constraints
We extend the evaluation to the following class of problems involving nonlinear constraints in both inputs and outputs, and perform high-dimensional analysis:
$\min_{y\in\mathbb{R}^{N_{O}}}f_{\text{obj}}(y)=\frac{1}{2}y^{T}Qy+p^{T}\sin{y},\quad\text{s.t.}\quad y^{T}Ay+c^{T}y+d=x^{3}.$ (10)
Here, $A\in\mathbb{R}^{N_{C}\times N_{O}\times N_{O}}$ denotes a tensor holding $N_{C}$ randomly generated symmetric matrices, while the remaining parameters follow the same sampling procedure as previously described. The varying parameter $x$ is uniformly drawn from the range $[-5,5]$, and the dataset dimensionality and split remain unchanged. We consider a problem with 200 variables and 150 nonlinear constraints.
The results are reported in Table 3. ENFORCE successfully predicts optimal solutions satisfying the set of nonlinear constraints across the whole test set. ENFORCE consistently achieves a 25× speedup in inference compared to the nonlinear programming solver, while maintaining the optimal objective within $6\%$. Traditional deep learning and soft constraint methods perform poorly when faced with nonlinear constraints, resulting in significant infeasibility or failure to approximate an optimal solution. Note that a comparison with DC3 is not included, as the original implementation does not support large-scale problem benchmarking involving nonlinear constraints. Once again, a scalability analysis across multiple problem dimensions is reported in Appendix C (Table 5).
Table 3: Performance comparison on nonconvex problems with nonlinear equality constraints involving 200 variables and 150 constraints. ENFORCE predicts feasible and near-optimal solutions across the entire test set, achieving a 25×25$\times$25 × speedup over IPOPT. Traditional deep learning and soft-constraint methods exhibit significant constraint violations and suboptimal performance.
Method Obj. value Max eq. Mean eq. Inf. time [s]
IPOPT [31] $-29.45\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $3.40\pm 1.40$
MLP $-53.07\pm 0.00$ $497.38\pm 4.64$ $118.39\pm 0.07$ $0.002\pm 0.001$
Soft ($\lambda_{c}=5$) $(3.02\pm 0.04)\times 10^{5}$ $79.23\pm 3.81$ $16.72\pm 0.07$ $0.001\pm 0.001$
Soft ($\lambda_{c}=1$) $(6.03\pm 0.04)\times 10^{4}$ $79.30\pm 3.70$ $16.68\pm 0.06$ $0.001\pm 0.000$
Soft ($\lambda_{c}=0.1$) $(6.11\pm 0.01)\times 10^{3}$ $78.35\pm 1.95$ $16.63\pm 0.01$ $0.002\pm 0.001$
ENFORCE $-27.77\pm 0.02$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.14\pm 0.08$
6 Conclusions
We propose ENFORCE, a method to ensure that neural network predictions satisfy a set of $\mathcal{C}^{1}$ nonlinear constraints $c(x,y)=0$ within a specified tolerance $\epsilon$, without relying on external solvers or incurring significant computational overhead. We prove the stability of the proposed projection mapping during training and provide theoretical insights into its convergence properties and applicability. The effectiveness of the method is demonstrated heuristically on (i) an illustrative function fitting case study selected for its interpretability and ease of visualization, and (ii) large-scale nonconvex optimization problems with nonlinear constraints. Our findings show that (1) ENFORCE consistently achieves constraint feasibility up to specified tolerance in the case studies, (2) task performance improves (up to 40%) when using ENFORCE over baseline methods in both function fitting and learning-based optimization settings, and (3) ENFORCE accelerates (×$\times 25$) computationally intensive tasks such as constrained optimization.
Acknowledgments and Disclosure of Funding
@@ 6 Conclusions
Appendix A Additional related works
Hard-constrained neural networks belong to the broader category of techniques incorporating inductive biases in deep learning models. Inductive biases guide the learning process toward compliance with domain knowledge or restrictions [40]. Architectures such as convolutional neural networks (CNNs) [41] and graph neural networks (GNNs) [42, 43] encode inductive biases by guaranteeing invariance with respect to patterns and symmetries. Simple analytical constraints can be enforced using differentiable functions, such as sigmoids or ReLU for output bounding and softmax for simplex constraints. Recent literature includes significant contributions for enforcing analytical inequality constraints, such as convex polytopes and convex sets more generally [44, 7, 45, 46, 47]. One can also constrain the neural network to guarantee specific functional characteristics, such as Lipschitz continuity [48] or Lyapunov stability [49].
Other methods have been proposed for constrained learning in NNs with analytical and logical expressions, mostly considering affine or convex regions [50, 51]. Many of them consider constraints only dependent on the input of the NN [52, 46, 34, 53], others design strategies to include the dependence on both inputs and outputs [47, 5]. Recently, contributions to enforce general logic and linear constraints have been proposed by the neuro-symbolic AI community, developing loss terms or constraining layers using logic programming [54, 55, 56]. For a broad and recent review on hard-constrained NNs, the reader is also referred to [16].
Appendix B Mathematical derivations
This section provides the key mathematical derivations, theorems, and proofs underlying the proposed method. These derivations are intended to illustrate the theoretical foundations of the approach and to support some of the discussions presented in the main text.
B.1 Closed-form neural projection
@@
B.2 Deviation from Newton’s method
To support the discussion raised in Section 4.1, we show how our method deviates from standard Newton’s method for solving nonlinear KKT conditions. Given a nonlinear program:
@@ B.2 Deviation from Newton’s method
Linearizing the system according to Newton’s iteration at $y=y_{0}$ results in: